OpenStudy (anonymous):

Let h(x) = 4 - 3 x^3. Compute h'( 3 ). Use this to find the equation of the tangent line to the curve y = 4 - 3 x^3 at the point ( 3 , -77 ). The equation of this tangent line can be written in the form y = mx+b. Determine m and b.

hartnn (hartnn):

could u calculate h'(x) ?

hartnn (hartnn):

h'(x) = -6(x)^2 so h'(3)= - 6(3)^2 = -6*9 = -54 that is the slope of your line m= -54 u have a point (3,-77) can u find the equation of line ?

OpenStudy (anonymous):

wait... how'd you get -6(x)^2?

hartnn (hartnn):

what is the derivative of x^n ?

hartnn (hartnn):

u started with derivatives ?

OpenStudy (anonymous):

... idk what that is.. my professor taught this today.. but she has a strong accent and im so confused.

hartnn (hartnn):

ok, \(\huge \frac{d}{dx}x^n=n.x^{n-1}\) remember this formula. According to this can u tell me derivative of x^3 ??

hartnn (hartnn):

is that ^^ making any sense ?

OpenStudy (anonymous):

I never was told that formula.

hartnn (hartnn):

so now u know..... can u tell me \(\frac{d}{dx}x^3\)